Find the Inverse Transform L {F(s)} of F(s) 38+2 $2 -38 +2 o L 1{F(s)} = 8e2t - 5et OL-{F(s)} 5 8-1 $ 2 L{F(s)} = -5e2-8et
Answer #1

3st2 f(3) = 5 - 35+2 F(S) = 35+2 3-25-5+2 Fis) 3572 S(S-2)-1(5-2) 3572 FIS) (5-2)(S-1) Using Portal fraction, B A + 35+2 (5-2t i {F(5y= at f(t) =18 - 5